# Models for rational $(\infty,1)$-categories

### v1

@@ 4. The model category for rational spaces
Proposition 4.2 ([Hir03, Proposition 1.1.11]).
If $A$ and $X$ are simplicial sets and $X$ is a Kan complex, then there is a natural weak homotopy equivalence $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(A,X)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|A|,|X|)$.
@@
Lemma 4.3 ([Hir03, Corollary 9.3.3(2)]).
If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of CW complexes and $Z$ is a topological space, then $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f,Z)\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$ is a weak homotopy equivalence.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(\operatorname{Sing}(S^{n}),K)$ is a weak homotopy equivalence for every $n\geq 2$.
@@
    The map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|\operatorname{Sing}(S^{n})|,|K|)$ is a weak homotopy equivalence for every $n\geq 2$.
@@
    The map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},|K|)$ is a weak homotopy equivalence for every $n\geq 2$.
@@
    a weak equivalence if and only if, for every $M$-local Kan complex $K$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(f,K)$ is a weak homotopy equivalence; and
@@ 5.1. Simplicial spaces
In the Reedy model structure on $\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}$, the weak equivalences are the levelwise weak homotopy equivalences, the cofibrations are the monomorphisms, and the fibrations are described using coskeleta; see [Ree, Theorem A] and [Ber18, Theorems 2.6.6 and 2.6.10]. This model structure is combinatorial, left proper, and simplicial; in particular, given two simplicial spaces $X$ and $Y$, we define their function complex $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(X,Y)$ by
@@
A property of the Reedy model structure on simplicial spaces that is relevant to our work is that it is cartesian [Ber18, Example 2.9.4]. In particular, given two simplicial spaces $X$ and $Y$, we have a simplicial space $Y^{X}$, and if $Y$ is Reedy fibrant, then so is $Y^{X}$. Although the localization of a cartesian model category need not be cartesian, we have a convenient criterion for localizations of the Reedy model structure on simplicial spaces.
Proposition 5.5 ([Rez01, Proposition 9.2]).
@@ 5.2. Complete Segal spaces
We gather some background on complete Segal spaces from [Ber18, Chapter 5].
@@
Definition 5.6 ([Rez01, §4.1]).
A simplicial space $W$ is a Segal space if $W$ is Reedy fibrant and, for every $n\geq 2$, its $n$-th Segal map is a weak homotopy equivalence.
@@
As the Segal maps are weak homotopy equivalences, we can define a composition operation on mapping spaces that is well-defined, associative, and unital up to a notion of homotopy. This notion of homotopy gives rise to the homotopy category of a Segal space. A Dwyer-Kan equivalence is a map that preserves this structure.
Definition 5.8 ([DK80, §2.4]).
@@
    for all $x,y\in\operatorname{ob}(X)$, the map $\operatorname{map}_{X}(x,y)\rightarrow\operatorname{map}_{Y}(f(x),f(y))$ is a weak homotopy equivalence; and
@@
    the induced functor on homotopy categories $\operatorname{Ho}(X)\rightarrow\operatorname{Ho}(Y)$ is an equivalence of categories.
@@
Definition 5.9 ([Rez01, §6]).
A Segal space $W$ is complete if the map $W_{0}\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(E^{t},W)$ is a weak homotopy equivalence.
@@
    a weak equivalence if and only if, for every complete Segal space $W$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(f,W)$ is a weak homotopy equivalence; and
@@
Lastly, we can complete a Segal space up to Dwyer-Kan equivalence in a functorial fashion as follows.
Theorem 5.12 ([Rez01, §14]).
@@ 6. Segal categories
We review another model for $(\infty,1)$-categories, that of Segal categories. Like complete Segal spaces, Segal categories are simplicial spaces. However, in place of the completeness condition, we ask that the space of 0-simplices of a Segal category be discrete. We recall one of Bergner’s model categories for Segal categories, and we show that it can be left-induced from Rezk’s model category for complete Segal spaces. We also go over Bergner’s proof of the equivalence between complete Segal spaces and Segal categories.
@@ 6.1. Definition of a Segal category
We first gather the definitions that we need from [Ber18, Chapter 6.1].
@@
A Segal category is a Segal precategory $X$ such that, for every $n\geq 2$, its $n$-th Segal map is a weak homotopy equivalence.
A Segal category is not required to be Reedy fibrant, so it need not be a Segal space. Still, objects and mapping spaces are defined as in Segal spaces, and the discreteness of the space of 0-simplices has a key consequence.
@@
If $X$ is a Segal category, then its space of $1$-simplices decomposes as $X_{1}=\coprod\limits_{(x,y)}\operatorname{map}_{X}(x,y)$. Thus, for every $n\geq 2$, the $n$-th Segal map of $X$ is the weak homotopy equivalence
@@
6.2. A model category for Segal categories
Unlike the Reedy model structure on simplicial spaces, there exists no model structure on Segal precategories with levelwise weak equivalences and levelwise cofibrations. Thus, the model category for Segal categories that we rely on does not arise as a left Bousfield localization. In [Ber07c], Bergner directly develops two model categories for Segal categories. The weak equivalences are described by means of the following result, which refers to the model category for complete Segal spaces $\mathcal{CSS}$.
Proposition 6.4 ([Ber07c, §5]).
There is a functor $L_{c}\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}\rightarrow\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}$ such that, for every Segal precategory $X$, its image $L_{c}(X)$ is a Reedy fibrant Segal category that is weakly equivalent to $X$ in $\mathcal{CSS}$.
Proposition 6.4 has the following consequence.
@@ 6.2. A model category for Segal categories
Let $\psi$ be a map of Segal precategories. Then, $\psi$ is a weak equivalence in $\mathcal{CSS}$ if and only if $L_{c}(\psi)$ is a weak equivalence in $\mathcal{CSS}$.
@@
Lemma 6.6 ([Ber07c, Lemma 5.9]).
If a map of Segal precategories $\psi$ has the right lifting property against the class of monomorphisms of Segal precategories, then $L_{c}(\psi)$ is a weak equivalence in $\mathcal{CSS}$.
We are now ready to describe the model category for Segal categories that we use in this paper.
Theorem 6.7 ([Ber07c, Theorem 5.1 and Corollary 5.13]; [Ber07a, Theorem 3.2]; [Ber17, Proposition 6.3]; [Pel, Theorem 6.4.4]).
@@
    a weak equivalence if and only if $L_{c}(f)$ is a weak equivalence in $\mathcal{CSS}$; and
@@
Lemma 6.5 now amounts to the following result.
Proposition 6.8.
Let $\psi$ be a map of Segal precategories. Then, $\psi$ is a weak equivalence in $\mathcal{CSS}$ if and only if $\psi$ is a weak equivalence in $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$.
@@ 6.3. The equivalence with complete Segal spaces
Showing that both complete Segal spaces and Segal categories are models for $(\infty,1)$-categories amounts to showing that the model categories $\mathcal{CSS}$ and $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ are Quillen equivalent. In [Ber07c], Bergner gives a Quillen equivalence by means of a right adjoint to the inclusion $I\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightarrow\mathcal{CSS}$ that sends a simplicial space $W$ to the Segal precategory $R(W)$ defined as the pullback
@@ 6.4. An approach by way of left transfer
We end this section by showing that the adjunction from Theorem 6.9 not only establishes an equivalence between complete Segal spaces and Segal categories, but also allows us to left-induce the model category $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ from the model category $\mathcal{CSS}$ in the sense of the following theorem.
Theorem 6.11 ([BHK+15, Theorem 2.23]).
Let $F\colon\mathcal{D}\rightleftarrows\mathcal{M}\colon G$ be an adjunction, where $\mathcal{D}$ is a locally presentable category and $\mathcal{M}$ is a combinatorial model category with class of cofibrations $C$. Suppose that, if a morphism $\psi$ in $\mathcal{D}$ has the right lifting property against the class of morphisms $F^{-1}(C)$, then $F(\psi)$ is a weak equivalence in $\mathcal{M}$. Then, there is a combinatorial model structure on $\mathcal{D}$ in which a map $\psi$ is
@@
In the setting of Theorem 6.11, we say that the model structure on $\mathcal{D}$ is a left-induced model structure.
@@
Firstly, we verify that the condition in Theorem 6.11 holds: if a map $\psi$ of Segal precategories has the right lifting property against the class of monomorphisms of Segal precategories, then $L_{c}(\psi)$ is a weak equivalence in $\mathcal{CSS}$ by Lemma 6.6, and therefore $\psi$ is a weak equivalence in $\mathcal{CSS}$ by Lemma 6.5. The resulting left-induced model structure on the category $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}$ of Segal precategories has the same cofibrations as $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$, as well as the same weak equivalences by Proposition 6.8. Hence, the left-induced model structure on $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}$ is $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. ∎
@@ 7. Rational Segal categories
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We introduce $M$-local Segal categories, which model $(\infty,1)$-categories enriched in $M$-local spaces. In particular, we introduce rational Segal categories, which model rational $(\infty,1)$-categories. We left Bousfield localize Bergner’s model category for Segal categories to get a model category whose fibrant objects are the Reedy fibrant $M$-local Segal categories. Thus, we produce a model category whose fibrant objects are the Reedy fibrant rational Segal categories.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A Segal category $X$ is $M$-local if, for any two objects $x$ and $y$ of $X$, their mapping space $\operatorname{map}_{X}(x,y)$ is $M$-local. A Segal category $X$ is rational if, for any two objects $x$ and $y$ of $X$, their mapping space $\operatorname{map}_{X}(x,y)$ is rational.
Proposition 6.3 implies that a Segal category $X$ is $M$-local if and only if $X$ is levelwise $M$-local, and we can encode the latter condition in the language of function complexes using our toolkit from Section 4. To that end, we use the fact that the inclusion of Segal precategories into simplicial spaces has a left adjoint that takes in a simplicial space $X$ and collapses the space of 0-simplices to its components to output a Segal precategory $X_{r}$.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a Reedy fibrant Segal category. The following statements are equivalent.
@@
    The Reedy fibrant Segal category $X$ is levelwise $M$-local.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X_{i})\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(\operatorname{Sing}(S^{n}),X_{i})$ is a weak homotopy equivalence for every $i\geq 0$ and every $n\geq 2$.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(\Delta[i]^{t}\times\operatorname{Sing}(S^{n}),X)$ is a weak homotopy equivalence for every $i\geq 0$ and every $n\geq 2$.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r},X)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r},X)$ is a weak homotopy equivalence for every $i\geq 0$ and every $n\geq 2$.
Proof.
Firstly, (2) and (3) are equivalent by Proposition 4.4. Then, (3) and (4) are equivalent by Corollary 5.4. Lastly, (4) and (5) are equivalent by adjointness. ∎
Proposition 7.2 instructs us to localize Bergner’s model category for Segal categories $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ at the set
$\widetilde{T}_{M^{-1}\mathbb{Z}}=\{(\Delta[i]^{t}\times\operatorname{Sing}(S^{n}))_{r}\rightarrow(\Delta[i]^{t}\times\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}))_{r}\mid i\geq 0\textrm{ and }n\geq 2\}.$
Theorem 7.3.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a combinatorial, left proper, and simplicial model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ on the category of Segal precategories in which a map $f$ is
  1. (1)
    a weak equivalence if and only if, for every Reedy fibrant $M$-local Segal category $X$, the induced map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(f,X)$ is a weak homotopy equivalence; and
@@
A Segal precategory $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ if and only if $X$ is a Reedy fibrant $M$-local Segal category.

### v2

@@ 3.3. The model category for rational spaces
Proposition 3.6 ([18, Proposition 1.1.11]).
If $A$ and $X$ are simplicial sets and $X$ is a Kan complex, then there is a natural weak equivalence $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(A,X)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|A|,|X|)$.
@@
Lemma 3.7 ([18, Corollary 9.3.3(2)]).
If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of CW complexes and $Z$ is a topological space, then $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f,Z)\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$ is a weak equivalence.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(\operatorname{Sing}(S^{n}),K)$ is a weak equivalence for all $n\geq 2$.
@@
    The map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|\operatorname{Sing}(S^{n})|,|K|)$ is a weak equivalence for all $n\geq 2$.
@@
    The map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},|K|)$ is a weak equivalence for all $n\geq 2$.
@@
    a weak equivalence if and only if, for every $M$-local Kan complex $K$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(f,K)$ is a weak equivalence; and
@@ 4.1. Simplicial spaces
In the Reedy model structure on $\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}$, the weak equivalences are the levelwise weak equivalences, the cofibrations are the monomorphisms, and the fibrations are described using coskeleta; see [24, Theorem A] and [7, Theorems 2.6.6 and 2.6.10]. This model structure is combinatorial, left proper, and simplicial; in particular, given two simplicial spaces $X$ and $Y$, we define their function complex $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(X,Y)$ by
@@
A property of the Reedy model structure on simplicial spaces that is relevant to our work is that it is cartesian [7, Example 2.9.4]. In particular, given two simplicial spaces $X$ and $Y$, we have a simplicial space $Y^{X}$, and if $Y$ is Reedy fibrant, then so is $Y^{X}$. Although the localization of a cartesian model category need not be cartesian, we have a convenient criterion for localizations of the Reedy model structure on simplicial spaces.
Proposition 4.5 ([25, Proposition 9.2]).
@@ 4.2. Complete Segal spaces
We give an overview of complete Segal spaces from [7, Chapter 5] and [25].
@@
Definition 4.6 ([25, §4.1]).
A simplicial space $W$ is a Segal space if $W$ is Reedy fibrant and, for every $n\geq 2$, its $n$-th Segal map is a weak equivalence.
@@
As the Segal maps are weak equivalences, we can define a composition operation on mapping spaces that is well-defined, associative, and unital up to a notion of homotopy. This notion of homotopy gives rise to the homotopy category of a Segal space. A Dwyer-Kan equivalence is a map that preserves this structure; the name refers to Dwyer and Kan’s analogous notion of equivalence for simplicial categories from [11, §2.4].
Definition 4.8 ([25, §7.4]).
@@
    for all $x,y\in\operatorname{ob}(X)$, the map $\operatorname{map}_{X}(x,y)\rightarrow\operatorname{map}_{Y}(f(x),f(y))$ is a weak equivalence; and
@@
    the induced functor on homotopy categories $\operatorname{Ho}(X)\rightarrow\operatorname{Ho}(Y)$ is essentially surjective.
@@
Definition 4.9 ([25, §6]).
A Segal space $W$ is complete if $W_{0}\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(E^{t},W)$ is a weak equivalence.
@@
    a weak equivalence if and only if, for every complete Segal space $W$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(f,W)$ is a weak equivalence; and
@@
Moreover, we can complete a Segal space up to Dwyer-Kan equivalence in a functorial fashion as follows.
Theorem 4.12 ([25, §14]).
@@
This result helps establish a categorical interpretation of the weak equivalences in $\mathcal{CSS}$ between Segal spaces.
Theorem 4.13 ([25, Theorem 7.7]).
A map of Segal spaces is a weak equivalence in $\mathcal{CSS}$ if and only if it is a Dwyer-Kan equivalence.
@@ 5. Segal categories
We review another model for $(\infty,1)$-categories, that of Segal categories. Segal categories are simplicial spaces; however, in place of the completeness condition, we ask that the space of 0-simplices of a Segal category be discrete. We recall Bergner’s model categories for Segal categories, and we show that Bergner’s model category for Reedy fibrant Segal categories can be left-induced from Rezk’s model category for complete Segal spaces. We also go over Bergner’s proof of the equivalence between complete Segal spaces and Segal categories.
@@ 5.1. Definition of a Segal category
We first gather the definitions that we need from [7, Chapter 6.1].
@@
A Segal category is a Segal precategory $X$ such that, for every $n\geq 2$, its $n$-th Segal map is a weak equivalence.
A Segal category is not required to be Reedy fibrant, so it need not be a Segal space. Still, objects and mapping spaces are defined as in Segal spaces. The discreteness of the space of 0-simplices has a key consequence.
@@
If $X$ is a Segal category, then its space of $1$-simplices decomposes as $X_{1}=\coprod\limits_{(x,y)}\operatorname{map}_{X}(x,y)$. Thus, for every $n\geq 2$, the $n$-th Segal map of $X$ is the weak equivalence
@@
5.2. Model categories for Segal categories
Unlike the Reedy model structure on simplicial spaces, there exists no model structure on Segal precategories with levelwise weak equivalences and levelwise cofibrations. Thus, the model categories for Segal categories that we rely on do not arise as left Bousfield localizations. In [5], Bergner directly develops two model categories for Segal categories. The weak equivalences are described by means of the following result, which refers to the model category for complete Segal spaces $\mathcal{CSS}$.
Proposition 5.4 ([5, §5]).
There is a functor $L_{c}\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}\rightarrow\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}$ such that, for every Segal precategory $X$, its image $L_{c}(X)$ is a Segal category and a Segal space that is weakly equivalent to $X$ in $\mathcal{CSS}$.
Definition 5.5 ([5, §5]).
A map of Segal precategories $f$ is a Dwyer-Kan equivalence if $L_{c}(f)$ is a Dwyer-Kan equivalence of Segal spaces.
In this terminology, Proposition 5.4 has the following consequence.
@@ 5.2. Model categories for Segal categories
Let $\psi$ be a map of Segal precategories. Then, $\psi$ is a weak equivalence in $\mathcal{CSS}$ if and only if $\psi$ is a Dwyer-Kan equivalence.
@@
Lemma 5.7 ([5, Lemma 5.9]).
If a map of Segal precategories $\psi$ has the right lifting property against the class of monomorphisms of Segal precategories, then $\psi$ is a Dwyer-Kan equivalence.
We describe the model category for Segal categories that we mainly use; the simplicial structure is established in [6, Proposition 6.3], and the fibrant objects are described in [5, Corollary 5.13] and [3, Theorem 3.2].
Theorem 5.8 ([5, Theorem 5.1]; [22, Theorem 6.4.4]).
@@
    a weak equivalence if and only if $f$ is a Dwyer-Kan equivalence; and
@@
Bergner also develops a Quillen equivalent model category for levelwise fibrant Segal categories; its left properness is proven in [7, Proposition 6.5.6], and its fibrant objects are described in [3, Theorem 4.2].
Theorem 5.9 ([5, Theorems 7.1 and 7.5]).
There is a combinatorial and left proper model structure $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{f}$ on the category of Segal precategories in which
  1. (1)
    a map $f$ is a weak equivalence if and only if $f$ is a Dwyer-Kan equivalence; and
  2. (2)
    the fibrant objects are the levelwise fibrant Segal categories.
Moreover, the adjunction $\operatorname{id}\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{f}\rightleftarrows\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\colon\operatorname{id}$ is a Quillen equivalence.
@@ 5.3. The equivalence with complete Segal spaces
In [5], Bergner gives a Quillen equivalence between complete Segal spaces and Segal categories by means of a right adjoint $R$ to the inclusion $I\colon\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}\rightarrow\mathcal{CSS}$. This right adjoint sends a simplicial space $W$ to the Segal precategory $R(W)$ defined as the pullback
@@ 5.4. An approach by way of left transfer
We show that the adjunction from Theorem 5.10 allows us to left-induce the model category $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ from the model category $\mathcal{CSS}$ in the sense of the following theorem.
Theorem 5.12 ([2, Theorem 2.23]).
Let $F\colon\mathcal{D}\rightleftarrows\mathcal{M}\colon G$ be an adjunction, where $\mathcal{D}$ is a locally presentable category and $\mathcal{M}$ is a combinatorial model category with class of cofibrations $C$. Suppose that, if a morphism $\psi$ in $\mathcal{D}$ has the right lifting property against the class of maps $F^{-1}(C)$, then $F(\psi)$ is a weak equivalence. Then, there is a combinatorial model structure on $\mathcal{D}$, called the left-induced model structure, in which a map $\psi$ is
@@
Firstly, we verify that the condition in Theorem 5.12 holds. If a map $\psi$ of Segal precategories has the right lifting property against the class of monomorphisms of Segal precategories, then $\psi$ is a Dwyer-Kan equivalence by Lemma 5.7, and therefore $\psi$ is a weak equivalence in $\mathcal{CSS}$ by Lemma 5.6. The resulting left-induced model structure on the category $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}$ of Segal precategories has the same cofibrations as $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$, as well as the same weak equivalences by Lemma 5.6. Hence, the left-induced model structure on $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}$ is $\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$. ∎
@@ 6. Rational Segal categories
Let $T$ be a set of simplicial maps. We introduce $T$-local Segal categories, which model $(\infty,1)$-categories enriched in $T$-local spaces. In particular, we introduce rational Segal categories, which model rational $(\infty,1)$-categories. We localize Bergner’s model category for Reedy fibrant Segal categories to get a model category whose fibrant objects are the Reedy fibrant $T$-local Segal categories. Thus, we produce a model category whose fibrant objects are the Reedy fibrant rational Segal categories. The same localization yields a Quillen equivalent model category whose fibrant objects are the levelwise fibrant $T$-local Segal categories. In particular, we get a Quillen equivalent model category whose fibrant objects are the levelwise fibrant rational Segal categories.
@@
Let $T$ be a set of simplicial maps. A Segal category $X$ is $T$-local if, for any two objects $x$ and $y$ of $X$, their mapping space $\operatorname{map}_{X}(x,y)$ is $T$-local. A Segal category $X$ is rational if, for any two objects $x$ and $y$ of $X$, their mapping space $\operatorname{map}_{X}(x,y)$ is rational.
Proposition 5.3 implies that a Segal category $X$ is $T$-local if and only if $X$ is levelwise $T$-local, and we can encode the latter condition in the language of function complexes. To that end, we use the fact that the inclusion of Segal precategories into simplicial spaces has a left adjoint that takes in a simplicial space $X$ and collapses the space of 0-simplices to its path components to output a Segal precategory $X_{r}$. Since Reedy fibrancy implies levelwise fibrancy [18, Corollary 15.3.12(2)], we formulate our characterization under levelwise fibrancy.
@@
Let $T$ be a set of simplicial maps. Let $X$ be a levelwise fibrant Segal category. The following statements are equivalent.
@@
    For every $i\geq 0$ and every $\lambda$ in $T$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(\lambda,X_{i})$ is a weak equivalence.
@@
    For every $i\geq 0$ and every $\lambda$ in $T$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(\Delta[i]^{t}\times\lambda,X)$ is a weak equivalence.
@@
    For every $i\geq 0$ and every $\lambda$ in $T$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}((\Delta[i]^{t}\times\lambda)_{r},X)$ is a weak equivalence.
Proof.
Firstly, (1) and (2) are equivalent by Corollary 4.4. Then, (2) and (3) are equivalent by adjointness. ∎
Proposition 6.2 instructs us to localize Bergner’s model categories for Segal categories at the set
$\widetilde{T}=\{(\Delta[i]^{t}\times\lambda)_{r}\mid i\geq 0\textrm{ and }\lambda\in T\}.$
Theorem 6.3.
Let $T$ be a set of simplicial maps. There is a combinatorial, left proper, and simplicial model structure $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ on the category of Segal precategories in which a map $f$ is
@@
    a weak equivalence if and only if, for every Reedy fibrant $T$-local Segal category $X$, the induced map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}^{\Delta^{\mathrm{op}}}}(f,X)$ is a weak equivalence; and
@@
A Segal precategory $X$ is fibrant in $\mathcal{L}_{T}\mathcal{S}\mathrm{e}\mathcal{C}\mathrm{at}_{c}$ if and only if $X$ is a Reedy fibrant $T$-local Segal category.
@@
Remark 6.5.
Localizing Bergner’s model category for levelwise fibrant Segal categories at the same set $\widetilde{T}$ yields a Quillen equivalent model category whose fibrant objects are the levelwise fibrant $T$-local Segal categories. In particular, there is a model category whose fibrant objects are the levelwise fibrant rational Segal categories, and it is Quillen equivalent to the model category for Reedy fibrant rational Segal categories.